OpenStudy (anonymous):

The volume of a sphere is found using the formula V=4/3 πr^3. The area of a circle is found using the formula A= πr^2 . What is the ratio of the volume to the area?

OpenStudy (campbell_st):

well write the formulas in ratio form \[\frac{4}{3} \pi r^3 : \pi r^2\] divide both by the common factor... and what do you get..?